# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v1

@@ B.1 Semantic Anomalies
The first relates to the distinction between low-level sensory anomalies and high-level semantic anomalies. Semantic anomalies might be present because a logical rule or a specific semantic constraint is violated. In contrast, AnomalyDINO is developed with low-level sensory anomalies in mind.
@@
Consider, for instance, the anomalies shown in Figure 5. While the cable with anomaly type ‘Bent Wire’ (Figure 5(b)) contains patches that cannot be well matched to patches of the reference image, this does not apply to the anomaly ‘Cable Swap’ (Figure 5(c)). The latter shows a semantic anomaly with two blue wires, while a nominal image of a ‘Cable’ should depict all three different cable types. As a result, all test patches of Figure 5(b) can be matched well with reference patches in $\mathcal{M}$, and thus, AnomalyDINO does not detect this anomaly type. Evaluating the detection performance for this specific failure case, ‘Cable Swap’, our proposed method essentially performs on chance level, giving a detection AUROC of 50.2% ($\pm$ 4.9%).
@@ B.2 The importance of informative reference samples
Figure 6: Example of an uninformative reference sample in category ‘Capsule’ (MVTec-AD). Some reference samples do not resemble the full concept of normality (here, the sample in Figure 6(b) does not show the text on the capsule such that any nominal sample with text will get a high anomaly score). Anomaly maps predicted by AnomalyDINO based on the depicted reference sample.
@@
(b) Anomaly map predicted by AnomalyDINO (1-shot) without (left) and with masking (right).
Figure 7: Visualization of the effect of masking in the presence of high background noise for the category ‘Capsules’ in VisA. As in Figure 4 we depict the anomaly map for ‘Capsules’ using a different colormap (red instead of yellow) for better visibility. Best viewed on a higher zoom level.
@@ C.1 Preprocessing
As discussed in Section 3, we consider two potential preprocessing steps in our pipeline: masking and rotations. We mask out irrelevant background patches, whenever the zero-shot segmentation of DINOv2 captures the object correctly (see Figure 2). Discarding background patches helps to mitigate the problem of potential background noise, thereby reducing the number of false positives. We find that some samples of anomalous samples also show artifacts in the background. An example of the category ‘Capsules’ from VisA is depicted in Figure 7. Without masking, the method would correctly predict the depicted sample to show an anomaly, but for the wrong reason (high anomaly scores for the background noise). We conclude that masking is essential to faithfully detect anomalies whenever higher variations in the background are expected. A quantitative analysis is provided in the next paragraphs.
In addition to masking (which decreases the size of $\mathcal{M}$), we consider augmenting the reference sample with rotations (which increases the size of $\mathcal{M}$). Here we distinguish the ‘agnostic’ scenario, where we do not know a priori about potential rotations, and augment by default. Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the test time.
The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in the test set (to comply with the one-shot setting), are given in Table 7.
Table 7: Default preprocessing steps for MVTec-AD and VisA. We do not mask textures, indicated by (T). In addition, we do not apply masking when the masking test failed, indicated by (MT). See Section 3 and Figure 2 for further discussion and a visualization.
@@
As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ influence the detection performance. The right preprocessing is in some cases more important than others. The full results are given in Figures 8 and 9.
Figure 8: Effect of preprocessing on MVTec-AD. Anomaly detection of AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
Figure 9: Effect of preprocessing on VisA. Anomaly detection with AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
@@
Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance. Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 90.1% in the 1-shot setting. This is intuitive, as the test samples of ‘Screw’ are taken from various angles. With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
Figure 10: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD). The left anomaly map is estimated from a reference sample with rotations (‘agnostic’), while the right anomaly map is based on a reference sample without rotations (‘informed’).
The same holds—although to a lesser extent—for the categories ‘Hazelnut’, and ‘Metal nut’ (both MVTec-AD), as well as for ‘Macaroni2’, ‘Cashew’, ‘PCB2’ and ‘PCB4’ categories (all VisA). However, we also observe that rotations of the reference sample can actually decrease the detection performance in some categories, namely ‘PCB1’ and ‘PCB3’ (both VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 4, right column). And rotations of nominal samples may falsely reduce the distance between those patches depicting bent connectors and the nominal memory bank $\mathcal{M}$. Such a failure case based on a suboptimal preprocessing decision is depicted in Figure 10. The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 10).
@@
Across the board, we see that masking improves the detection performance. The only exception is ‘Candle’ from VisA with a slight decrease in performance. On further inspection, we find that the estimated foreground masks of DINOv2 occasionally fail to capture all relevant patches (the decision on whether to mask or not is inferred from the first nominal sample only, 1-shot setting). We leave the development of more robust and adaptive masking techniques for future work.
@@
The design choices in our pipeline influence the test time. To see the potential effects we measure the test time per sample on MVTec for the different combinations of masking (yes/no) and rotation (yes/no) for the full pipeline. This includes preparation of the memory bank $\mathcal{M}$ (i.e., feature extraction and the optional augmentation and masking operations), as well as the inference time for all test samples (i.e., feature extraction and optional masking). The results are given in Figure 11. The runtime was assessed utilizing a single NVIDIA A40 GPU, consistently employed throughout all experiments detailed in this paper (each experiment can be executed on a single NVIDIA A40 GPU).
The resulting run time increases by a factor of 4, when comparing the baseline (no masking, no rotation) to the scenario where masking and rotations are applied to each product. Not rotating a reference sample in cases where it is unnecessary (the difference between ‘informed’ and ‘agnostic’) lowers the test time by approximately 30%.
Figure 11: Effect on masking and augmentation on test time (per sample) in the 1-shot setting with AnomalyDINO. We measure the test time on a single NVIDIA A40 GPU and report mean and standard deviation over three independent runs on MVTec-AD.
@@ C.2 Scoring
A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow Roth et al., 2022, and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$. We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
Table 8: Effect of aggregation statistics $q$ on the detection performance on MVTec-AD and VisA. $\mathcal{D}$ denotes the set of patch distances to the nominal memory bank, $H_{0.01}(\mathcal{D})$ the 1% highest entries thereof, and $\mathcal{A}$ denotes the anomaly map derived from $\mathcal{D}$ (anomaly scores for each pixel, after upsampling and smoothing, see Section 3). All results in %.
Scoring $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ $q=\mathrm{max}(\mathcal{\mathcal{A}})$ $q=\mathrm{max}(\mathcal{D})$
Shots AUROC F1-max AP AUROC F1-max AP AUROC F1-max AP
MVTec-AD 1 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 95.8$\pm$0.6 95.4$\pm$0.3 97.7$\pm$0.4 95.1$\pm$0.5 94.8$\pm$0.1 97.4$\pm$0.4
2 96.9$\pm$0.8 96.6$\pm$0.4 98.2$\pm$0.7 96.2$\pm$1.1 96.0$\pm$0.4 97.7$\pm$0.7 95.6$\pm$1.0 95.4$\pm$0.5 97.4$\pm$0.6
4 97.7$\pm$0.1 97.1$\pm$0.3 98.5$\pm$0.3 97.2$\pm$0.2 96.7$\pm$0.3 98.1$\pm$0.4 96.7$\pm$0.2 96.1$\pm$0.1 97.9$\pm$0.4
8 98.2$\pm$0.1 97.5$\pm$0.1 99.1$\pm$0.2 97.7$\pm$0.1 97.1$\pm$0.3 98.8$\pm$0.1 97.3$\pm$0.2 96.5$\pm$0.4 98.6$\pm$0.1
16 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 98.0$\pm$0.1 97.4$\pm$0.2 99.1$\pm$0.1 97.7$\pm$0.2 96.9$\pm$0.3 98.9$\pm$0.1
VisA 1 86.2$\pm$1.4 83.5$\pm$1.0 87.7$\pm$1.4 85.1$\pm$1.7 82.7$\pm$1.0 87.1$\pm$1.3 83.8$\pm$1.8 82.0$\pm$1.0 85.8$\pm$1.6
2 88.9$\pm$1.9 85.3$\pm$1.4 90.4$\pm$1.4 87.7$\pm$1.9 84.3$\pm$1.4 89.7$\pm$1.2 86.6$\pm$1.9 83.4$\pm$1.2 88.5$\pm$1.2
4 91.5$\pm$0.7 87.6$\pm$0.9 92.6$\pm$0.6 90.7$\pm$0.3 86.4$\pm$0.6 92.1$\pm$0.3 89.8$\pm$0.3 85.7$\pm$0.5 91.2$\pm$0.5
8 92.9$\pm$0.1 89.0$\pm$0.3 93.8$\pm$0.1 92.0$\pm$0.1 87.8$\pm$0.3 93.2$\pm$0.3 91.2$\pm$0.1 87.0$\pm$0.2 92.4$\pm$0.4
16 94.0$\pm$0.0 90.0$\pm$0.1 94.8$\pm$0.1 93.3$\pm$0.2 88.9$\pm$0.4 94.4$\pm$0.2 92.6$\pm$0.3 88.2$\pm$0.4 93.7$\pm$0.3
The results in Table 8 show that just the maximum patch distance leads to already good results, but upsampling and smoothing the patch distances improves upon that (maximum of $\mathcal{A}$). Taking the mean of all patch distances above the 99% quantile ($q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$) improves above the standard choice. We did not optimize over the percentile and instead fixed it to 99%. Typically, the number of patches per image ranges between 200 and 1000 (depending on resolution and masking), such that 2 between 10 patches are considered. For specific products/objects and expected anomaly types, other statistics $q$ might be (more) suitable.

### v2

@@ B.1 Semantic Anomalies
The first relates to the distinction between low-level sensory anomalies and high-level semantic anomalies. Semantic anomalies might be present because a logical rule or a specific semantic constraint is violated. In contrast, AnomalyDINO is developed with low-level sensory anomalies in mind. Consider, for instance, the anomalies shown in Figure 6.
@@
While the cable with anomaly type ‘Bent Wire’ (Figure 6(b)) contains patches that cannot be well matched to patches of the reference image, this does not apply to the anomaly ‘Cable Swap’ (Figure 6(c)). The latter shows a semantic anomaly with two blue wires, while a nominal image of a ‘Cable’ should depict all three different cable types. As a result, all test patches of Figure 6(b) can be matched well with reference patches in $\mathcal{M}$, and thus, AnomalyDINO does not detect this anomaly type. Evaluating the detection performance for this specific failure case, ‘Cable Swap’, our proposed method essentially performs on chance level, giving a detection AUROC of 50.2% ($\pm$ 4.9%).
@@ B.2 The importance of informative reference samples
Figure 7: Example of an uninformative reference sample in category ‘Capsule’ (MVTec-AD). Some reference samples do not resemble the full concept of normality (here, the sample in Figure 7(b) does not show the text on the capsule, i.e., any nominal sample with text be considered anomalous). Anomaly maps predicted by AnomalyDINO based on the depicted reference sample.
@@
(b) Anomaly map predicted by AnomalyDINO (1-shot) without masking (left) and with masking (right).
Figure 8: Visualization of the effect of masking in the presence of high background noise for the category ‘Capsules’ in VisA (1-shot). As in Figure 5 we depict the anomaly map for ‘Capsules’ using a different colormap (red instead of yellow) for better visibility. Best viewed on a higher zoom level.
@@ C.1 Preprocessing
As discussed in Section 3, we consider two potential preprocessing steps in our pipeline: masking and rotations. We mask out irrelevant background patches, whenever the zero-shot segmentation of DINOv2 captures the object correctly (see Figure 2). Discarding background patches helps to mitigate the problem of potential background noise, thereby reducing the number of false positives. A representative example, here from the category ‘Capsules’ from the VisA dataset, is depicted in Figure 8. Without masking, the method would correctly predict the depicted sample to show an anomaly, but for the wrong reason (high anomaly scores caused by background noise/contamination). In contrast, the irrelevant background areas are discarded by our masking approach, such that the anomalous region is correctly identified. We conclude that masking is essential to faithfully detect anomalies whenever higher variations in the background are expected—and as the issue of low variations is particularly pressing in the few-shot regime, caused by the minimal amount of nominal reference samples, suitable masking can significantly boost the performance in these cases. This is showcased by the superior anomaly localization performance A quantitative analysis is provided in the next paragraphs.
In addition to masking (which decreases the size of $\mathcal{M}$), we consider augmenting the reference sample with rotations (which increases the size and diversity of $\mathcal{M}$). Here we distinguish the ‘agnostic’ scenario, where we do not know a priori about potential rotations, and augment by default. Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the size of the memory bank $\mathcal{M}$, the time to construct $\mathcal{M}$, as well as the test time itself.
Table 8: Default preprocessing steps for MVTec-AD and VisA. We do not mask textures, as indicated by (T). In addition, we do not apply masking when the masking test on the first train sample failed, as indicated by (MT). See Section 3 and Figure 2 for further discussion and visualization.
@@
Figure 9: Effect of preprocessing on MVTec-AD. Anomaly detection of AnomalyDINO-S (448) in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in $X_{\mathrm{ref}}$ (to comply with the one-shot setting), are given in Table 8.
Figure 10: Effect of preprocessing on VisA. Anomaly detection with AnomalyDINO-S (448) in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs). For VisA, the ‘informed’ scenario is equivalent to only applying masking (all categories), while ‘agnostic’ is equivalent to masking and augmentations (all categories), see Table 8.
@@
As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ and preprocess test samples, influence the detection performance. The results per object for MVTec-AD and VisA are given in Figures 9 and 10, respectively.
@@
Figure 11: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD). The left anomaly map is estimated from a reference sample with rotations (‘agnostic’), and the anomaly is not detected. In contrast, the right anomaly map is based on a reference sample without rotations (‘informed’), and the anomaly is successfully detected. This example highlights the importance of a carefully designed preprocessing pipeline for each object.
Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance. Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 89.2% in the 1-shot setting. This is intuitive, as the test samples of ‘Screw’ are taken from various angles. With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
The same holds—although to a lesser extent—for the categories ‘Hazelnut’, ’Cable’ and ‘Wood’ (all MVTec-AD). However, we also observe that rotations of the reference sample can also decrease the detection performance in some categories, namely ‘PCB1/2/3’ and ‘Macaroni1’ (all VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 5, right column). And rotations of nominal samples may falsely reduce the distance between those patches depicting bent connectors and the nominal memory bank $\mathcal{M}$. Such a failure case based on a suboptimal preprocessing decision is depicted in Figure 11. The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 11). This highlights the importance of carefully designing the pre- and postprocessing pipeline for each object/product considered.
@@
Figure 12: Masking examples from MVTec-AD for all categories that passed the masking test (on a single train sample, see Table 8). Examples have not been cherry-picked.
We utilize the zero-shot masking capabilities of DINOv2 to keep the overhead for this preprocessing step minimal. By applying a threshold to the first principal component [29], we can typically distinguish between the background and the foreground.
We also employ a straightforward rule-based strategy to enhance the robustness and generalizability of the PAC-based masking in industrial settings: ensuring that the center crop of the image is predominantly occupied by the object of interest. This minor adjustment is necessary because industrial images often differ significantly from the data on which DINOv2 was originally trained. In addition, we use dilation and morphological closing to improve the quality of the mask. In our experiments, we applied masking only to the test samples as the size of $\mathcal{M}$ did not matter in the few-shot regime. Across the board, we see that the proposed masking technique improves the detection performance—in many categories even significantly, e.g., for ‘Capsules’, ‘Macaroni1/2’, or ‘PCB1/2/3/4’ (all VisA). We leave further improvements and the exploration of more advanced masking techniques for future work.
Figure 13: Masking examples from VisA (not cherry-picked).
Table 9: Runtime analysis on MVTec-AD (mean and std of inference time over all 1725 test samples, and mean and std of time to populate the memory bank for each object) for different shots (Tab. 9(a)), preprocessing choices (Tab. 9(b)), sample resolutions (Tab. 9(c)) and model sizes (Tab. 9(d)). All times are reported in seconds, measured on a single NVIDIA A40 with GPU warmup and CUDA kernel synchronization. The default setting is 1-shot, agnostic preprocessing, model size S, and resolution 448 (underlined for reference).
Shots Inference Memory Bank
1 0.060$\pm$0.012 00.52$\pm$0.040
2 0.059$\pm$0.012 00.85$\pm$0.070
4 0.059$\pm$0.012 01.58$\pm$0.080
8 0.060$\pm$0.011 03.05$\pm$0.170
16 0.063$\pm$0.012 06.02$\pm$0.390
full (masking only) 0.067$\pm$0.010 16.62$\pm$4.430
full (agnostic) 0.130$\pm$0.044 93.72$\pm$20.56
(a) Runtime in dependence of shots.
Mask? Rotate? Inference Memory Bank
no no 0.055$\pm$0.011 0.17$\pm$0.03
no yes 0.055$\pm$0.012 0.51$\pm$0.05
yes no 0.068$\pm$0.012 0.18$\pm$0.03
yes yes 0.067$\pm$0.012 0.51$\pm$0.04
informed 0.059$\pm$0.016 0.22$\pm$0.11
agnostic 0.060$\pm$0.012 0.52$\pm$0.04
(b) Runtime in dependence of preprocessing choices.
Resolution Inference Memory Bank
224 0.043$\pm$0.010 0.31$\pm$0.04
448 0.060$\pm$0.012 0.52$\pm$0.04
672 0.086$\pm$0.014 0.75$\pm$0.06
896 0.141$\pm$0.021 1.26$\pm$0.05
(c) Runtime in dependence of image resolution.
Model Size Inference Memory Bank
S (21 M) 0.060$\pm$0.012 0.52$\pm$0.04
B (86 M) 0.084$\pm$0.021 0.77$\pm$0.05
L (300 M) 0.141$\pm$0.029 1.24$\pm$0.06
G (1,100 M) 0.306$\pm$0.034 2.67$\pm$0.14
(d) Runtime in dependence of model size.
@@
The design choices in our pipeline influence the run time of the proposed method. To see the potential effects we measure the inference time per sample on MVTec-AD in various scenarios together with the time to set up the memory bank $\mathcal{M}$. The runtime was assessed utilizing a single NVIDIA A40 GPU, consistently employed throughout all experiments detailed in this paper (each experiment can be executed on a single NVIDIA A40 GPU). The runtime is measured with GPU warmup and CUDA kernel synchronization for a fair comparison.
The results are given in Table 9. The average inference time per sample with AnomalyDINO-S (448) amounts to approximately 60ms in the few-shot regime ($\approx$ 16.7 samples/s), and only moderately increases with a larger memory bank (to 67ms ($+11\%$) for the full-shot scenario without augmentations). Compared to SOTA competitors in the one-shot regime, this is at least one order of magnitude faster (see Figure 3). Without any preprocessing steps, the 1-shot inference time amounts to approximately 55ms per sample ($\approx$ 18 samples/s). When applying both masking and rotations, the runtime increases from 55ms to 67ms, an moderate increase of roughly 23% (compared to the scenario without any preprocessing steps). The informed scenario can reduce the time to build the memory bank, but inference time is only affected for larger sample sizes. As expected, higher resolutions and larger architectures lead to increased runtimes.
@@ C.2 Scoring
Table 10: Effect of aggregation statistics $q$ on the detection performance on MVTec-AD and VisA, evaluated for AnomalyDINO-S (448). $\mathcal{D}$ denotes the set of patch distances to the nominal memory bank, $H_{0.01}(\mathcal{D})$ the 1% highest entries thereof, and $\mathcal{A}$ denotes the anomaly map derived from $\mathcal{D}$ (anomaly scores for each pixel, after upsampling and smoothing, see Section 3). All results in %.
Scoring $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ $q=\mathrm{max}(\mathcal{\mathcal{D}})$ $q=\mathrm{max}(\mathcal{A})$
Shots AUROC F1-max AP AUROC F1-max AP AUROC F1-max AP
MVTec-AD 1 96.5$\pm$0.4 96.0$\pm$0.2 98.1$\pm$0.3 95.0$\pm$0.5 94.7$\pm$0.2 97.4$\pm$0.4 94.9$\pm$0.7 94.6$\pm$0.6 97.5$\pm$0.4
2 96.7$\pm$0.8 96.5$\pm$0.4 98.1$\pm$0.7 95.5$\pm$1.0 95.3$\pm$0.5 97.4$\pm$0.6 95.0$\pm$1.3 94.8$\pm$0.8 97.4$\pm$0.9
4 97.6$\pm$0.1 97.0$\pm$0.3 98.4$\pm$0.3 96.6$\pm$0.2 96.0$\pm$0.1 97.8$\pm$0.4 96.3$\pm$0.4 95.7$\pm$0.2 98.1$\pm$0.3
8 98.0$\pm$0.1 97.4$\pm$0.1 99.0$\pm$0.2 97.2$\pm$0.1 96.5$\pm$0.4 98.6$\pm$0.1 97.0$\pm$0.0 96.3$\pm$0.1 98.6$\pm$0.1
16 98.3$\pm$0.1 97.7$\pm$0.2 99.3$\pm$0.0 97.6$\pm$0.2 96.9$\pm$0.3 98.9$\pm$0.1 97.4$\pm$0.1 96.6$\pm$0.2 98.8$\pm$0.1
VisA 1 85.6$\pm$1.5 83.1$\pm$1.1 86.6$\pm$1.3 82.4$\pm$1.9 81.4$\pm$1.0 84.0$\pm$1.7 80.5$\pm$1.3 80.6$\pm$0.8 82.1$\pm$0.9
2 88.3$\pm$1.8 84.8$\pm$1.2 89.2$\pm$1.3 85.1$\pm$1.8 82.5$\pm$1.3 86.4$\pm$1.1 83.3$\pm$1.8 82.2$\pm$1.1 84.5$\pm$1.5
4 91.3$\pm$0.8 87.5$\pm$1.0 91.8$\pm$0.7 88.4$\pm$0.3 84.9$\pm$0.6 89.0$\pm$0.4 86.8$\pm$1.4 84.3$\pm$0.9 87.6$\pm$1.3
8 92.6$\pm$0.1 88.6$\pm$0.2 92.9$\pm$0.2 90.0$\pm$0.3 86.3$\pm$0.2 90.6$\pm$0.5 88.8$\pm$0.5 85.3$\pm$0.3 89.7$\pm$0.3
16 93.8$\pm$0.1 89.9$\pm$0.3 94.2$\pm$0.3 91.6$\pm$0.5 87.2$\pm$0.7 92.2$\pm$0.4 90.5$\pm$0.3 86.8$\pm$0.5 91.5$\pm$0.4
@@
A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow [33] and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$. We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
The results in Table 10 show that just the maximum patch distance leads to already good results while upsampling and smoothing the patch distances giving slightly weaker results (maximum of $\mathcal{A}$).
Taking the mean of all patch distances above the 99% quantile ($q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$) improves above the standard choice. We did not optimize over the percentile and instead fixed it to 99%. Typically, the number of patches per image ranges between 200 and 1000 (depending on resolution and masking), such that 2 between 10 patches are considered. For specific products/objects and expected anomaly types, other statistics $q$ might be (more) suitable.